The base of an isosceles triangle is 8 m and the legs are 6m. What's is the hight?
Answer by edjones(8007)

The altitude divides the triangle into 2 right triangles with c=6 and b=4
a^2+b^2=c2
a^2+4^2=6^2
a^2+16=36
a^2=20
a=sqrt(20)
a=sqrt(5*4)
=2sqrt(5)m
.
Ed